OpenStudy (anonymous):

I have about 3 questions that need answer pretty quick. Please Help!

OpenStudy (anonymous):

1. Calculate the surface area of this triangular prism. a. 255 m^2 b. 360 m^2 c. 468 m^2 d. 495 m^2

OpenStudy (anonymous):

Did u mean Total Surface Area?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

First Try to find the base of the right angled triangle

OpenStudy (anonymous):

Is the formula SA =2B +(s1+s2+s3)h

OpenStudy (anonymous):

Use pythagorean theorem

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

 then, 9^2 +x^2 =15^2

OpenStudy (anonymous):

caCan't you just use the fact that the 12m on the bottom is parallel to the leg on the top?

OpenStudy (anonymous):

Oh yes! u can use it

OpenStudy (anonymous):

I still don't get it.

OpenStudy (anonymous):

Now, u can directly use the formula mentioned above

OpenStudy (anonymous):

But I don't know what the base is.

OpenStudy (anonymous):

It is 12

OpenStudy (anonymous):

So, Find the Area of the Base

OpenStudy (anonymous):

Okay. how do we do that?

OpenStudy (anonymous):

Area of traingle =1/2 base * height

OpenStudy (anonymous):

so 6 * 10.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Then what is the height?

OpenStudy (anonymous):

Base of the triangle is 12 and height is 9

OpenStudy (anonymous):

so 6 * 9 = 54

OpenStudy (anonymous):

54 is our base now?

OpenStudy (anonymous):

Yes, thats the Area of the right angled triangle base

OpenStudy (anonymous):

So SA =2(54) +(s1+s2+s3)h

OpenStudy (anonymous):

Now, Total Surface Area = 2* Area of base +(s1+s2+s3) * height of the prism

OpenStudy (anonymous):

The height of the prism is 10.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

SA =2(54) +(s1+s2+s3)10

OpenStudy (anonymous):

How do we find the sides?

OpenStudy (anonymous):

The sides are 9 , 12 and 15

OpenStudy (anonymous):

Observe the figure

OpenStudy (anonymous):

So, SA =2(54) +(9+12+15)10

OpenStudy (anonymous):

very correct

OpenStudy (anonymous):

The answer is 468